OpenStudy (anonymous):

which of the following are a posible solution to the (3x-1)(5x+20)=0 a.1/3,4 b.1/3,-4 c.1,20 d.-1,-20

myininaya (myininaya):

\[3x-1=0 \text{ or } 5x+20=0\]

OpenStudy (anonymous):

b

myininaya (myininaya):

We need to solve both of these.

OpenStudy (anonymous):

1/3 n 4

myininaya (myininaya):

x=1/3 or x=-4

OpenStudy (anonymous):

ohh transposing ..soo negative..its" b" srry